Giant​ Enterprises' stock has a required return of 15.1​%. The​ company, which plans to pay a dividend of $1.52 per share in the coming​ year, anticipates that its future dividends will increase at an annual rate consistent with that experienced over 2013​-2019 period, when the following dividends were​ paid:

Year 2019 2018 2017 Dividend per Share $1.46 $1.41 $1.35 $1.30 $1.25 $1.20 $1.16 2016 2015 2014 2013
a.If the​ risk-free rate is 7%,what is the risk premium on​ Giant's stock? round to one decimal place
b.Using the​ constant-growth model, estimate the value of​ Giant's stock.  
​(​Hint:
  Round the computed dividend growth rate to the nearest whole​ percent.)
c.  Explain what​ effect, if​ any, a decrease in the risk premium would have on the value of​ Giant's stock.
Answer #1

a). Risk premium on Giant's Stock = Required Return on Giant's Stock - Risk-free Rate

= 15.1% - 7% = 8.1%

b). g = [D(2019) / D(2013)][1/(2019 - 2013)] - 1

= [$1.46 / $1.16]1/6 - 1

= [1.2586]1/6 - 1

= 1.0391 - 1 = 0.0391, or 3.91%, or 4%

P(2019) = D(2020) / [r - g] = $1.52 / [0.151 - 0.04] = $1.52 / 0.111 = $13.69

c). If risk premium decreases, this means required return will be less, which will result into the increase in the value of Giant's stock.